OpenStudy (anonymous):

How would i go about interagting (x^2+1)^3*x

OpenStudy (anonymous):

Use a u substitution. In particular, letting \(u = x^2+1\) will work well here.

OpenStudy (anonymous):

a question: i found a similar problem at i am confused over this step: =\[1/5\int\limits_{}^{} u^4du\] then it goes to: =\[(1/5)(1/5)u^5+c\] Where does the other 1/5 come from?

OpenStudy (anonymous):

It comes from performing the integration.. \[\int x^a dx = \frac{x^{a+1}}{a+1} + C\]